

********************************************************************
UARS WEEKLY STATUS REPORT - 30 JUN 2000
********************************************************************
***********
SOLSTICE
***********
We continue to gather enough data to meet all the SOLSTICE science
objectives during this period of reduced tape recorder operations.
SOLSTICE operates normally, collecting solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible.
We have planned a special observation of Mercury for execution on Friday
(June 30). Additional observations will be made later in the year when
more observation time is available. Thanks to Rich Beck for his help.
Level 2 processing is halted due to bugs in the OA_SSPP_ATT and
OA_SAT_ATT services related to data gaps. CDHF and CSC personnel have
released engineering patches to these routines and we are beginning to
incorporate them into a test release of our production software.
***********
SUSIM
***********
SUSIM continues to gather a near normal set of daily solar scans
scheduled for periods of predicted telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Processing of instrument Daily Activity Plans continues with no
problems. We are continuing to coordinated with the Flight
Operations Team and to assist the Project Scientist in the
scheduling of instrument operations.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
- All requested reprocessing is complete as of this reporting period
CDHF REHOST:
- Successful processing on Rehost System:
Attitude Extraction, Observatory Log, Sensor Extraction, HALOE, HRDI,
MLS, SUSIM, Solstice, and WINDII.
- Waiting validation results from the HRDI group on successful
production testing done on the Rehost System.
- System test is progress.
***********
FOT
***********
REPORTING PERIOD
This report covers 17 June 2000 (Orbit 47912, GMT Day 169) through 23 June 2000
(Orbit 48017, GMT Day 175).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
WINDII 170/0139 47928 Power ON
WINDII 170/0721 47932 Power OFF
WINDII 170/2023 47939 Power ON
WINDII 171/0324 47944 Power OFF
The beta angle ranged between the angles of 31.0 degrees to 31.3 degrees for
this report period. The beta angle is now decreasing toward a minimum angle of
2.0 degrees on 04 July 2000 (DOY 186).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.0 to 1.3 degrees C. Additional battery performance data for each
battery on 16 June 2000 and 23 June 2000 are provided in Appendix D.
The clock error ranged from -11.3 to 11.1 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% OUT-OF-LIMITS = 03
% ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 23
% WINDII = 04
% HALOE = 02
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 01
Four (4) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
- Transponder A: 00/145 - 2287.497008
- Transponder B: 00/152 - 2287.498920
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/174 show the observatory data loss to be 296 hours, 20
minutes, 7.713 seconds (an increase of 0 minutes 0.0 seconds since the last
report period). This is a 0.40 percent data loss which equals a 99.60 percent
data capture for the mission.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT will finally continue testing the new TMONs again next week. The
coding changes lost in the UTTS disk crash have been reinserted and the
scenario lost is in the process of being rebuilt. Once the TMONs have been
thoroughly tested, they will be presented for review and approval before
uplink.
The FOT has begun writing the online procedure for safehold should it happen
before the new TMONs are in place or during and OBC halt. This procedure will
be tested on the simulator and put online for immediate execution should an
anomalous condition take place.
MEETINGS
The FOT held a teleconference with Dave Lorenz and the FSW group to discuss the
Multi-TDRS patch. Many options were discussed and the FOT and FSW group
continue to look into the best method for implementing this patch. A decision
for 3 or 4 TDRS will be given at the next FSW CCB.
The FOT attended a meeting to review the TMON trip on day 159. Lessons learned
and future best practices were discussed to avoid this problem in the future.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for July 04, 2000 (DOY 186). This maneuver
will be a Reverse to Forward Flight maneuver (Southern Hemisphere viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 23 June 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
47914 169/0320 00-278 NBTR PLAYBACK
47914 169/0356 00-278 NBTR NOOPS D-169
47923 169/1742 00-279 NBTR NOOPS D-170
47928 170/0139 00-280 WINDII Power ON
47930 170/0502 00-279 NBTR NOOPS D-170
47930 170/0540 00-279 NBTR PLAYBACK
47932 170/0721 00-281 WINDII Power OFF - S/C LOW POWER
47938 170/1744 00-282 NBTR NOOPS D-171
47939 170/2023 00-280 WINDII Power ON
47944 171/0324 00-281 WINDII Power OFF - S/C LOW PWR
47944 171/0324 00-282 NBTR PLAYBACK
47945 171/0504 00-282 NBTR NOOPS D-171
47953 171/1747 00-283 NBTR NOOPS D-172
47959 172/0326 00-283 NBTR PLAYBACK
47960 172/0506 00-283 NBTR NOOPS D-172
47968 172/1847 00-284 NBTR NOOPS D-173
47969 172/1929 00-285 OBC Clk Adj 0.00 to 20.6 msec/day
47974 173/0328 00-284 NBTR PLAYBACK
47975 173/0508 00-284 NBTR NOOPS D-173
47982 173/1609 00-284 NBTR NOOPS D-174
47989 174/0329 00-286 NBTR PLAYBACK
47990 174/0509 00-287 OBC Clk Adj 20.6 to 0.00 msec/day
47990 174/0549 00-286 NBTR NOOPS D-174
47995 174/12:43 00-289 HALOE FOV test
47995 174/1347 00-290 HALOE Reset FOV
47998 174/1758 00-288 NBTR NOOPS D-175
48004 175/0331 00-288 NBTR PLAYBACK
48006 175/0648 00-288 NBTR NBTR NOOPS D-175
48013 175/1953 00-292 NBTR NBTR NOOPS D-176
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 23 June 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPSEQBIMON Y-HI PSE QB Current Mon 47914-17
HRPTSPBIMON R-LO PTS PB Current Mon 47916-7
HRPSEQBIMON Y-HI PSE QB Current Mon 47930
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
-None this report period
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 23 June 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
47969 172/19:29 (20 June 00) from 0.00 to 20.6 msec/day
47990 174/05:09 (22 June 00) from 20.6 to 0.00 msec/day
_______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 23 June 2000
16 June 2000 - GMT Day 168 Beta = 31.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- ---------- ---------- ------ ------- ------
1 +39.2/+11.2 -2.42/-3.68 +0.0/+0.0 15.7 0.000 0.0
2 +28.0/-28.0 +7.09/+5.15 +28.8/-15.2 23.8 1.030 23.0
3 +22.4/-78.4 +5.47/+4.19 +28.4/-15.2 23.7 1.028 22.0
23 June 2000 - GMT Day 175 Beta = 31.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- ---------- ---------- ------ ------- ------
1 +39.2/+11.2 -3.05/-4.31 +0.0/+0.0 15.7 0.000 0.0
2 +16.8/-56.0 +5.47/+4.51 +28.8/-14.4 25.1 1.040 20.7
3 +22.4/-28.0 +4.51/+3.23 +28.0/-14.4 25.1 1.037 20.5
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
______________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 23 June 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-027 FOT 47928 170/0215 22524 Late Acq./Data Loss
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-064 FOT 44135 10/8-9/99 N/A HRDI Cal Lamp Errors
99-071 FOT 44398 299/1212 21512 TPF Not Receiv 512
00-026 FOT see list June 00 14170 Generic Late Acq.
00-027 FOT 47928 170/0215 22524 Late Acq./Data Loss
SEVNTFW00/25